# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v2

@@ 5 Improving Label DP Training with Retraining (RT)
Motivated by our theoretical results in Section 4 which show that retraining (abbreviated as RT henceforth) can improve accuracy in the presence of label noise, we propose to apply our proposals in Section 1, viz., full RT and more importantly, consensus-based RT to improve label DP training (because it involves noisy labels). Note that this can be done on top of any label DP mechanism and that too at no additional privacy cost (both the predicted labels and originally provided noisy labels are private). We empirically evaluate full and consensus-based RT on three classification datasets (available on TensorFlow) trained with label DP. These include two vision datasets, namely CIFAR-10 and CIFAR-100, and one language dataset, namely AG News Subset (Zhang et al.,, 2015). All the empirical results are averaged over three different runs. We only provide important experimental details here; the other details can be found in Appendix G.
@@
CIFAR-10/100. We train a ResNet-18 model on CIFAR-10 and CIFAR-100 with label DP. Label DP training is done with the prior-based method of Ghazi et al., (2021) – specifically, Alg. 3 with two stages. Our training set consists of 45k examples and we assume access to a validation set with clean labels consisting of 5k examples which we use for deciding when to stop training, setting hyper-parameters, etc. [^4] In practice, we do not need full access to the validation set. Instead, the validation set can be stored by a secure agent which returns us a private version of the validation accuracy and this will not be too far off from the true validation accuracy when the validation set is large enough. For CIFAR-10 and CIFAR-100 with three different values of $\epsilon$, we list the test accuracies of the baseline (i.e., the method of Ghazi et al., (2021)), full RT and consensus-based RT in Tables 1 and 2, respectively. Notice that consensus-based RT is the clear winner. Also, for the three values of $\epsilon$ in Table 1 (CIFAR-10), the size of the consensus set (used in consensus-based RT) is $\sim 31\%$, $55\%$ and $76\%$, respectively, of the entire training set. The corresponding numbers for Table 2 (CIFAR-100) are $\sim 11\%$, $34\%$ and $56\%$, respectively. So for small $\epsilon$ (high label noise), consensus-based RT comprehensively outperforms full RT and baseline with a small fraction of the training set. Further, in Table 3, we list the accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted labels (which are the same as the given labels) over the consensus set for CIFAR-10 and CIFAR-100. To summarize, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than full dataset.
@@
Table 1: CIFAR-10. Test set accuracies (mean $\pm$ standard deviation). Consensus-based RT is better than full RT which is better than the baseline.
@@
Table 2: CIFAR-100. Test set accuracies (mean $\pm$ standard deviation). Overall, consensus-based RT is significantly better than full RT which is somewhat better than the baseline.
@@
One may wonder how more model parameters affect the results due to potential overfitting. So here we consider training ResNet-34 (which has more parameters than ResNet-18) on CIFAR-100 with label DP; the setup is exactly the same as our previous experiments with ResNet-18. Please see the results and discussion in Table 4. In summary, consensus-based RT is pretty effective even with ResNet-34 but the amount of improvement here is less compared to ResNet-18 due to overfitting. Additionally, one may wonder if retraining is useful when a label noise-robust technique is used during initial training, i.e., in the baseline (the method of Ghazi et al., (2021)). To that end, we once again train ResNet-34 on CIFAR-100 but with the popular noise-correcting technique of “forward correction” (Patrini et al.,, 2017) applied to the first stage of our baseline (it is not clear how to apply it to the second stage). [^5] Forward correction worked the best among other popular noise-robust techniques such as backward correction (Patrini et al.,, 2017) and noise-robust loss functions (Wang et al.,, 2019; Zhang and Sabuncu,, 2018). Please see the result and discussion for $\epsilon=5$ in Table 5. In summary, retraining can offer gains even after noise-robust training.
Table 4: CIFAR-100 w/ ResNet-34. Test set accuracies (mean $\pm$ standard deviation). Just like Table 2 (ResNet-18), consensus-based RT is the clear winner here. However, note that the performance and amount of improvement with ResNet-34 is worse than ResNet-18 due to more overfitting because of more parameters; this is expected.
@@
Table 5: Noise-robust “forward correction” (Patrini et al.,, 2017) applied to baseline (CIFAR-100 w/ ResNet-34). Test set accuracies (mean $\pm$ standard deviation). We see that retraining can yield improvement even after noise-robust training, although the amount of improvement is less compared to Table 4 (no noise-correction).
@@
AG News Subset (https://www.tensorflow.org/datasets/catalog/ag_news_subset). This is a news article classification dataset consisting of 4 categories – world, sports, business or sci/tech. We reserve 10% of the given training set for validation and we use the rest for training with label DP. Just like the CIFAR experiments, we assume that the validation set comes with clean labels. We use the small BERT model available in TensorFlow and the BERT English uncased preprocessor; links to both of these are in Appendix G. We pool the output of the BERT encoder, add a dropout layer with probability = 0.2, followed by a softmax layer. We fine-tune the full model. Here, label DP training is done with randomized response. We list the test accuracies of the baseline (i.e., randomized response), full RT and consensus-based RT in Table 6 for three different values of $\epsilon$. Even here, consensus-based RT is the clear winner. For the three values of $\epsilon$ in Table 6, the size of the consensus set (used in consensus-based RT) is $\sim 28\%$, $32\%$ and $38\%$, respectively, of the entire training set. So here, consensus-based RT appreciably outperforms full RT and baseline with less than two-fifths of the entire training set. Finally, in Table 7, we list the accuracies of predicted labels and given labels over the entire dataset and accuracies of predicted labels (= given labels) over the consensus set. Even here, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This explains why consensus-based RT performs the best, even though the consensus set is much smaller than the full dataset.
@@
Table 7: AG News Subset. Accuracies over the entire dataset and over the consensus set. Conclusions are the same as Table 3.
0.3 $53.20\pm 2.82$ $32.52\pm 2.05$ $\bm{61.81}\pm 2.66$
0.5 $66.78\pm 1.31$ $35.5\pm 0.14$ $\bm{76.48}\pm 0.93$
0.8 $79.98\pm 0.80$ $42.53\pm 0.13$ $\bm{89.59}\pm 0.43$
So in summary, consensus-based RT significantly improves model accuracy.
Additional Empirical Results in the Appendix. In Appendix H, we show that consensus-based RT also outperforms retraining on samples for which the model is the most confident; this is similar to self-training’s method of sample selection in the semi-supervised setting. In Appendix I, we show that RT is beneficial even without a clean validation set. Furthermore, going beyond label DP, we show that consensus-based RT is also beneficial in the presence of human annotation errors which can be thought of as “real” label noise in Appendix J.
@@ 6 Conclusion
In this work, we provided the first theoretical result showing retraining with hard labels can provably increase model accuracy in the presence of label noise. We also showed the efficacy of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) in improving label DP training at no extra privacy cost. We will conclude by discussing some future directions of work. Our theoretical results in this work focus on the full retraining scheme. Given that consensus-based retraining worked very well empirically, we would like to analyze it theoretically as a future work. Another potential extension of our work is to analyze retraining under non-uniform label noise models. We also hope to test our ideas on larger scale models and datasets in the future.
@@ Appendix A Problem Setting of Figure 1
The setting is exactly the same as the problem setting in Section 4 with ${\bm{{\mu}}}=\gamma\bm{e}_{1}$, $\bm{\Sigma}=\bm{I}-\bm{e}_{1}\bm{e}_{1}^{\sf T}$, $d=50$ and $p=0.4$. In the direction of ${\bm{{\mu}}}$, we have $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}=y(1+u)\gamma^{2}$. We consider balanced classes (i.e., $\pi_{+}=\pi_{-}=\frac{1}{2}$) and choose $u\sim{\sf Unif}[0,4]$, the uniform distribution over the interval $[0,4]$. In Figure 1(a), $\gamma^{2}=0.5$ (large separation) and in Figure 1(b), $\gamma^{2}=0.3$ (small separation). The number of training samples in each case is 300 and the retraining is done on the same training set on which the model is initially trained. The learned classifiers from vanilla training and retraining are the same as in Section 4 (i.e., eq. 4.3 and eq. 4.7, respectively). Finally, the test accuracy of both vanilla training and retraining in Figure 1(b) is $68\%$.
@@
We first prove the upper bound $\alpha_{0}(\bm{x})$ and then the lower bound $\widetilde{\alpha}_{0}(\bm{x})$.
B.1 Upper bound $\alpha_{0}(\mathbf{x})$
@@
We next use a classical lower bound on $\Phi$. Let $\Phi^{c}(t)=1-\Phi(t)=\Phi(-t)$. Using (Abramowitz and Stegun,, 1968, Equation (7.1.13)), we have for any $t>0$:
@@
B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
@@
We next recall the tail bound of normal distribution (Abramowitz and Stegun,, 1968, Equation (7.1.13)):
@@
By invoking (B.4) we obtain
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)=y)\geq 1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,,$
which completes the derivation of $\widetilde{\alpha}_{0}(\bm{x})$.
@@
Using Theorem 4.1 we have:
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\leq\operatorname{\mathbb{E}}[\alpha_{0}(\bm{x})]$
$=1-\frac{1}{2\sqrt{2\pi}}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]$
$\leq 1-\frac{1}{2\sqrt{2\pi}}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\bigg{|}\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right]\mathbb{P}\left(\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right)$
$\leq 1-\frac{1}{2\sqrt{2\pi}}\exp\left(-5(1+\sqrt{n}(1-2p))^{2}v^{2}\right)\mathbb{P}\left(\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right)\,,$ (C.1)
@@
given that ${\bm{z}}$ and $u$ are independent. Next note that $\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}$ is distributed as a Gaussian vector in a $(d-1)$-dimensional space, and therefore $\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}^{2}$ is $\chi^{2}$ distribution with $(d-1)$-degree of freedom. We next use the following result about the tail bound of $\chi^{2}$ distribution to control $\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}$.
Lemma C.1
(Ghosh,, 2021, Theorem 4) Suppose $X\sim\chi^{2}_{d}(\lambda)$. Then for $0<\eta<d+\lambda$, we have
$\mathbb{P}(X<d+\lambda-\eta)\leq\exp\left[-\frac{d\eta^{2}}{4(d+2\lambda)^{2}}\right]\,.$
As a corollary of Lemma C.1, by choosing $\eta=\lambda+d/2$, we obtain $\mathbb{P}(X<d/2)\leq e^{-d/16}$ for any $\lambda\geq 0$. Using this we have
$\mathbb{P}\left(\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}^{2}<d/2\right)\leq e^{-d/16}\,.$ (C.4)
@@
Combining (C.4), (C.5), (C.3) and (C.2) with the bound (C.1), we arrive at
${\sf acc}({\widehat{\bm{\theta}}}_{0})\leq 1-\frac{1}{4\sqrt{2\pi}}\exp\left(-160(1+\sqrt{n}(1-2p))^{2}\frac{\gamma^{4}}{\lambda_{\min}^{2}d}\right)\left(1-e^{-d/16}\right)\,.$
To derive the lower bound on ${\sf acc}({\widehat{\bm{\theta}}}_{0})$ note that by using Theorem 4.1 we have
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\geq\operatorname{\mathbb{E}}[\widetilde{\alpha}_{0}(\bm{x})]$
$=1-\frac{1}{2}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (C.6)
@@
where the inequality holds because $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\gamma^{2}$ due to which $|\langle\bm{x},{\bm{{\mu}}}\rangle|\geq\gamma^{2}$ (given that $u$ is a non-negative random variable) and because,
$\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}=\left\|\bm{\Sigma}{\bm{z}}\right\|_{\ell_{2}}^{2}\leq\lambda_{\max}^{2}\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2\lambda_{\max}^{2}d$
when event $\mathcal{E}_{2}$ happens. Further, $\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and so using tail bounds for $\chi^{2}_{d}$, see e.g., (Ghosh,, 2021, Theorem 3), we have $\mathbb{P}(\mathcal{E}_{2}^{c})\leq e^{-d/8}$. Putting things together, we obtain

### v3

@@ 5 Improving Label DP Training with Retraining (RT)
Motivated by our theoretical results in Section 4 which show that retraining (abbreviated as RT henceforth) can improve accuracy in the presence of label noise, we propose to apply our proposals in Section 1, viz., full RT and more importantly, consensus-based RT to improve local label DP training because it involves training with noisy labels. Note that this can be done on top of any label DP mechanism and that too at no additional privacy cost (both the predicted labels and originally provided noisy labels are private). Here we empirically evaluate full and consensus-based RT on four classification datasets (available on TensorFlow) trained with label DP. These include three vision datasets, namely CIFAR-10, CIFAR-100, and DomainNet (Peng et al., 2019), and one language dataset, namely AG News Subset (Zhang et al., 2015). Unless otherwise mentioned, all our empirical results here are with the standard cross-entropy loss. All the results are averaged over three runs. We only provide important experimental details here; the other details can be found in Appendix H.
@@
CIFAR-10/100. We train a ResNet-18 model (from scratch) on CIFAR-10 and CIFAR-100 with label DP. Label DP training is done with the prior-based method of Ghazi et al. (2021) – specifically, Algorithm 3 with two stages. Our training set consists of 45k examples and we assume access to a validation set with clean labels consisting of 5k examples which we use for deciding when to stop training, setting hyper-parameters, etc. [^7] In practice, we do not need access to the validation set. Instead, it can be stored by a secure agent which returns us a private version of the validation accuracy and this will not be too far off from the true validation accuracy when the validation set is large enough. We also show some results without a validation set in Appendix K. For CIFAR-10 and CIFAR-100 with three different values of $\epsilon$ (DP parameter; see Def. 3.1), we list the test accuracies of the baseline (i.e., the method of Ghazi et al. (2021)), full RT and consensus-based RT in Tables 1 and 2, respectively. Notice that consensus-based RT is the clear winner. Also, for the three values of $\epsilon$ in Table 1 (CIFAR-10), the size of the consensus set (used in consensus-based RT) is $\sim 31\%$, $55\%$ and $76\%$, respectively, of the entire training set. The corresponding numbers for Table 2 (CIFAR-100) are $\sim 11\%$, $34\%$ and $56\%$, respectively. So for small $\epsilon$ (high label noise), consensus-based RT comprehensively outperforms full RT and the baseline with a small fraction of the training set. Further, in Table 3, we list the accuracies of the predicted labels and the given labels over the entire (training) dataset and the accuracy of the predicted labels (which are the same as the given labels) over the consensus set for CIFAR-10 and CIFAR-100. To summarize, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than full dataset.
@@
Table 1: CIFAR-10. Test set accuracies (mean $\pm$ standard deviation). Consensus-based RT is better than full RT which is better than the baseline (initial training).
@@
Table 2: CIFAR-100. Test set accuracies (mean $\pm$ standard deviation). Overall, consensus-based RT is significantly better than full RT which is somewhat better than the baseline (initial training).
@@
One may wonder how an increase in the number of model parameters affects results because of potential overfitting to label noise. So next, we train ResNet-34 (which has nearly six times the number of parameters of ResNet-18) on CIFAR-100; the setup is exactly the same as our previous ResNet-18 experiments. Additionally, one may wonder if retraining is useful when a label noise-robust technique is used during initial training, i.e., on top of the method of Ghazi et al. (2021). To that end, we train ResNet-34 on CIFAR-100 with (i) the popular noise-correcting technique of forward correction (Patrini et al., 2017) applied to the first stage of Ghazi et al. (2021) (it is not clear how to apply it to the second stage), and (ii) the noise-robust symmetric cross-entropy (CE) loss function of Wang et al. (2019) with $\alpha=0.8,\beta=0.2$ used instead of the standard CE loss function. We show the results for ResNet-34 without any noise-robust method, forward correction used initially, and with the symmetric CE loss function in Tables 4, 5, and 6, respectively, for $\epsilon=\{4,5\}$; please see the detailed discussion in the captions. In summary, consensus-based RT is pretty effective even with overfitting by ResNet-34. Also, more importantly, consensus-based RT is beneficial even after noise-robust initial training.
Table 4: CIFAR-100 w/ ResNet-34. Test set accuracies (mean $\pm$ standard deviation). Just like Table 2 (ResNet-18), consensus-based RT is the clear winner here. But the performance in case of ResNet-34 is worse than ResNet-18 due to more overfitting to noise because of more parameters; this is expected.
@@
Table 5: CIFAR-100 w/ ResNet-34: noise-robust forward correction (Patrini et al., 2017) used initially (in the baseline). Test set accuracies (mean $\pm$ standard deviation). We see that retraining (especially, consensus-based RT) yields improvement even after noise-robust initial training, although the amount of improvement is less compared to Table 4 (no noise-correction).
@@
Table 6: CIFAR-100 w/ ResNet-34: noise-robust symmetric CE loss (Wang et al., 2019) used instead of standard CE loss. Test set accuracies (mean $\pm$ standard deviation). We see that retraining (especially, consensus-based RT) yields improvement even with a noise-robust loss function.
$\epsilon$ Baseline Full RT Consensus-based RT
4 $37.07\pm 2.03$ $38.17\pm 2.03$ $\mathbf{43.20}\pm 1.77$
5 $53.10\pm 0.54$ $53.40\pm 0.33$ $\mathbf{56.13}\pm 0.25$
DomainNet (https://www.tensorflow.org/datasets/catalog/domainnet). This is an image classification dataset much larger than CIFAR where the images belong to one of 345 classes. Here we do linear probing (i.e., fitting a softmax layer) on top of a ResNet-50 pretrained on ImageNet. We reserve 10% of the entire training set for validation and use the rest for training with label DP. Just like the CIFAR experiments, we assume that the validation set comes with clean labels. In Table 7, we show results for $\epsilon=\{3,4\}$ when label DP training is done with the method of Ghazi et al. (2021) with two stages (same as the CIFAR experiments). Observe that consensus-based RT is the clear winner. Further, in Tables 8 and 9, we show the corresponding results when the noise-robust techniques of forward and backward correction (Patrini et al., 2017) are applied to the first stage of Ghazi et al. (2021) in initial training (it is not clear how to apply forward and backward correction to the second stage). As expected, forward & backward correction lead to better initial model performance (compared to no correction). The main thing to note however is that consensus-based RT yields significant gains even after noise-robust initial training, consistent with our earlier results. [^8] It is worth noting that for $\epsilon=3$, consensus-based RT leads to similar accuracy with and without noise correction.
Table 7: DomainNet. Test set accuracies (mean $\pm$ standard deviation). Consensus-based RT leads to very significant improvement.
$\epsilon$ Baseline Full RT Consensus-based RT
3 $23.60\pm 0.92$ $29.23\pm 1.03$ $\mathbf{36.30}\pm 0.75$
4 $48.25\pm 0.05$ $52.10\pm 0.10$ $\mathbf{57.40}\pm 0.20$
Table 8: DomainNet: noise-robust forward correction used initially. Test set accuracies (mean $\pm$ standard deviation). Consensus-based RT yields significant improvement even after noise-robust initial training.
$\epsilon$ Baseline Full RT Consensus-based RT
3 $31.23\pm 0.56$ $33.30\pm 0.65$ $\mathbf{36.07}\pm 0.78$
4 $58.50\pm 0.08$ $58.63\pm 0.12$ $\mathbf{61.80}\pm 0.08$
Table 9: DomainNet: noise-robust backward correction used initially. Test set accuracies (mean $\pm$ standard deviation). Once again, consensus-based RT yields significant improvement even after noise-robust initial training.
$\epsilon$ Baseline Full RT Consensus-based RT
3 $30.17\pm 0.61$ $31.47\pm 0.74$ $\mathbf{35.03}\pm 0.78$
4 $56.63\pm 0.37$ $56.80\pm 0.37$ $\mathbf{60.47}\pm 0.46$
Finally, we show results on a language dataset.
@@
AG News Subset (https://www.tensorflow.org/datasets/catalog/ag_news_subset). This is a news article classification dataset consisting of 4 categories – world, sports, business or sci/tech. Just like the previous experiments, we keep 10% of the full training set for validation which we assume comes with clean labels, and use the rest for training with label DP. We use the small BERT model available in TensorFlow and the BERT English uncased preprocessor; links to both of these are in Appendix H. We pool the output of the BERT encoder, add a dropout layer with probability = 0.2, followed by a softmax layer. We fine-tune the full model. Here label DP training is done with randomized response. We list the test accuracies of the baseline (i.e., randomized response), full RT and consensus-based RT in Table 10 for three different values of $\epsilon$. Once again, consensus-based RT is the clear winner. For the three values of $\epsilon$ in Table 10, the size of the consensus set is $\sim 28\%$, $32\%$ and $38\%$, respectively, of the entire training set. So here, consensus-based RT appreciably outperforms full RT and baseline with less than two-fifths of the entire training set. In Table 11 (Appendix I), we list the accuracies of the predicted & given labels over the full dataset and the consensus set; the observations are similar to Table 3 giving us an idea of why consensus-based RT performs the best.
@@
So in summary, consensus-based retraining is a straightforward post-processing step that can be applied on top of a base method to significantly improve its performance in the presence of noisy labels.
Additional empirical results in the Appendix. In Appendix J, we show that consensus-based RT outperforms retraining on samples for which the model is the most confident; this is similar to self-training’s method of sample selection in the semi-supervised setting. In Appendix K, we show that RT is beneficial even without a validation set. Finally, going beyond label DP, we show that consensus-based RT is beneficial in the presence of human annotation errors which can be thought of as “real” label noise in Appendix L.
@@ 6 Conclusion
In this work, we provided the first theoretical result showing retraining with predicted hard labels can provably increase model accuracy in the presence of label noise. We also showed the efficacy of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given label) in improving local label DP training. We will conclude by discussing some limitations of our work which pave the way for future directions of work. Our theoretical results in this work focused on full retraining. Because consensus-based retraining worked very well empirically, we would like to analyze it theoretically in the future. Also, our theoretical results are under the uniform label noise model. In the future, we would like to analyze retraining under non-uniform label noise models. We also hope to test our ideas on larger scale models and datasets.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
Adel Javanmard is supported in part by the NSF Award DMS-2311024, the Sloan fellowship in Mathematics, an Adobe Faculty Research Award, and an Amazon Faculty Research Award. The authors are grateful to anonymous reviewers for their feedback on improving this paper.
@@ Appendix A Problem Setting of Figure 1
The setting is exactly the same as the problem setting in Section 4 with $d=50$, ${\bm{{\mu}}}=\gamma\bm{e}_{1}$, $\bm{\Sigma}=\bm{I}-\bm{e}_{1}\bm{e}_{1}^{\sf T}$, and $p=0.4$. In the direction of ${\bm{{\mu}}}$, we have $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}=y(1+u)\gamma^{2}$. We consider balanced classes (i.e., $\pi_{+}=\pi_{-}=\frac{1}{2}$) and choose $u\sim{\sf Unif}[0,4]$, the uniform distribution over the interval $[0,4]$. In Figure 1(a), $\gamma^{2}=0.5$ (large separation) and in Figure 1(b), $\gamma^{2}=0.3$ (small separation). The number of training samples in each case is 300 and the retraining is done on the same training set on which the model is initially trained. The learned classifiers from initial training and retraining are the same as in Section 4 (i.e., Equations (4.3) and (4.8), respectively). Finally, the test accuracy of both initial training and retraining in Figure 1(b) is $68\%$.
@@
We first prove the lower bound $\alpha_{0}(\bm{x})$ and then the upper bound $\widetilde{\alpha}_{0}(\bm{x})$.
B.1 Lower bound $\alpha_{0}(\mathbf{x})$
@@
We next use a classical lower bound on $\Phi$. Let $\Phi^{c}(t)=1-\Phi(t)=\Phi(-t)$. Using Equation (7.1.13) of Abramowitz & Stegun (1968), we have for any $t>0$:
@@
B.2 Upper bound $\widetilde{\alpha}_{0}(\mathbf{x})$
@@
We next recall the tail bound of normal distribution in Equation (7.1.13) of Abramowitz & Stegun (1968):
@@
By invoking (B.4), we obtain the desired bound $\widetilde{\alpha}_{0}(\bm{x})$.
@@
We will first derive the lower bound on ${\sf err}({\widehat{\bm{\theta}}}_{0})$. Using Theorem 4.1 (and because the training set $\mathcal{T}$ has measure zero), we have
${\sf err}({\widehat{\bm{\theta}}}_{0})$ $\geq\operatorname{\mathbb{E}}[\alpha_{0}(\bm{x})]$
$=\frac{1}{2\sqrt{2\pi}}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]$
$\geq\frac{1}{2\sqrt{2\pi}}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\bigg{|}\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right]\mathbb{P}\left(\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right)$
$\geq\frac{1}{2\sqrt{2\pi}}\exp\left(-5(1+\sqrt{n}(1-2p))^{2}v^{2}\right)\mathbb{P}\left(\frac{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\leq v^{2}\right)\,,$ (C.1)
@@
given that ${\bm{z}}$ and $u$ are independent. Next note that $\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}$ is distributed as a Gaussian vector in a $(d-1)$-dimensional space, and therefore $\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}^{2}$ follows the chi-squared distribution with $(d-1)$ degrees of freedom. Using the tail bound of the chi-squared distribution (see Theorem 4 of Ghosh (2021)), we get:
$\mathbb{P}\left(\left\|\mathcal{P}^{\perp}_{\bm{\mu}}\bm{z}\right\|_{\ell_{2}}^{2}<d/2\right)\leq e^{-d/16}\implies\mathbb{P}\left(\left\|\mathcal{P}^{\perp}_{\bm{\mu}}{\bm{z}}\right\|_{\ell_{2}}\geq\sqrt{d/2}\right)\geq 1-e^{-d/16}.$ (C.4)
@@
Combining (C.4), (C.5), (C.3) and (C.2) with (C.1) and plugging in our choice of $v=\frac{4\sqrt{2}\gamma^{2}}{\lambda_{\min}\sqrt{d}}$, we arrive at
${\sf err}({\widehat{\bm{\theta}}}_{0})\geq\frac{1}{4\sqrt{2\pi}}\exp\left(-160(1+\sqrt{n}(1-2p))^{2}\frac{\gamma^{4}}{\lambda_{\min}^{2}d}\right)\left(1-e^{-d/16}\right)\,.$
Let us now derive the upper bound on ${\sf err}({\widehat{\bm{\theta}}}_{0})$. Using Theorem 4.1 (and because the training set $\mathcal{T}$ has measure zero), we have
${\sf err}({\widehat{\bm{\theta}}}_{0})$ $\leq\operatorname{\mathbb{E}}[\widetilde{\alpha}_{0}(\bm{x})]$
$=\frac{1}{2}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]+2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (C.6)
@@
where the inequality holds because $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\gamma^{2}$ due to which $|\langle\bm{x},{\bm{{\mu}}}\rangle|\geq\gamma^{2}$ (given that $u$ is a non-negative random variable) and because
$\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}=\left\|\bm{\Sigma}{\bm{z}}\right\|_{\ell_{2}}^{2}\leq\lambda_{\max}^{2}\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2\lambda_{\max}^{2}d,$
when event $\mathcal{E}_{2}$ happens. Further, $\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and so using tail bounds for $\chi^{2}_{d}$ (see Theorem 3 of Ghosh (2021)), we have $\mathbb{P}(\mathcal{E}_{2}^{c})\leq e^{-d/8}$. Putting things together, we obtain